OpenStudy (anonymous):

Solve 11^5x = 104.  Round to the nearest ten-thousandth.

OpenStudy (anonymous):

This is a logarithm question. 11^5x = 104 Lg both sides, lg 11^5x = lg 104 Bring down the power of 5x ( you can do it when it is in the logarithm form) 5x lg 11= lg 104 5x= lg104/ lg 11 5x = 1.93686 x = 0.38372 round of to nearest ten-thousandth, 0.3837

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

Np.